Gaseous ethane will react with gaseous oxygen to produce gaseous carbon dioxide and gaseous water . Suppose 2.7 g of ethane is mixed with 3.63 g of oxygen. Calculate the maximum mass of carbon dioxide that could be produced by the chemical reaction. Be sure your answer has the correct number of significant digits.
Answer #1



Molar mass of C2H6,
MM = 2*MM(C) + 6*MM(H)
= 2*12.01 + 6*1.008
= 30.068 g/mol


mass(C2H6)= 2.7 g

use:
number of mol of C2H6,
n = mass of C2H6/molar mass of C2H6
=(2.7 g)/(30.07 g/mol)
= 8.98*10^-2 mol

Molar mass of O2 = 32 g/mol


mass(O2)= 3.63 g

use:
number of mol of O2,
n = mass of O2/molar mass of O2
=(3.63 g)/(32 g/mol)
= 0.1134 mol
Balanced chemical equation is:
2 C2H6 + 7 O2 ---> 4 CO2 + 6 H2O


2 mol of C2H6 reacts with 7 mol of O2
for 8.98*10^-2 mol of C2H6, 0.3143 mol of O2 is required
But we have 0.1134 mol of O2

so, O2 is limiting reagent
we will use O2 in further calculation


Molar mass of CO2,
MM = 1*MM(C) + 2*MM(O)
= 1*12.01 + 2*16.0
= 44.01 g/mol

According to balanced equation
mol of CO2 formed = (4/7)* moles of O2
= (4/7)*0.1134
= 6.482*10^-2 mol


use:
mass of CO2 = number of mol * molar mass
= 6.482*10^-2*44.01
= 2.853 g
Answer: 2.9 g